OpenStudy (anonymous):

What is the volume of a cylinder with the following dimensions Radius 6ft Height 24ft A. 33.14 ft2 B. 2712.96 ft2 C. 33.14 ft3 D. 2712.96 ft3

OpenStudy (nurali):

V=pi*r^2*h r=6 ft and h=24 ft

OpenStudy (nurali):

put in value and tell me answer.

OpenStudy (anonymous):

so the answer is c ?

OpenStudy (nurali):

No solve it

OpenStudy (nurali):

pi=3.14

OpenStudy (nurali):

r^2 means r*r

OpenStudy (anonymous):

i got A is dhat correct ?

OpenStudy (nurali):

Wrong

OpenStudy (nurali):

V=3.14*6*6*24 V=2712.96 ft3

OpenStudy (anonymous):

i really dont know the answer to this question so can you give me the answer

OpenStudy (nurali):

2712.96 ft3